OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (anonymous):

Is it together or are they two separate problems?

OpenStudy (anonymous):

They are together.

OpenStudy (phi):

3x + 4y = -2 2x - 3y = 10 one way is to multiply the top equation (that means every term BOTH sides) by -2 and multiply the 2nd equation by +3 what do you get for the 2 new equations ?

OpenStudy (anonymous):

Is the new equations -6x+-8y=4 and 6x+9y=30? Or did I do something wrong?

OpenStudy (phi):

close . redo the 2nd equation.

OpenStudy (anonymous):

Do you change the + to a -?

OpenStudy (phi):

3*(2x-3y)= 3*10 6x -9y = 30

OpenStudy (phi):

now add the 2 equations -6x+ -8y=4 6x+ -9y=30

OpenStudy (anonymous):

Like this? 6x-9y=30 or is it like this? 6x-(-9y)=30?

OpenStudy (phi):

are you asking what you get when you multiply the 2nd equation by 3 ? 2x + -3y = 10 3 times each term: 3*2x + 3* -3y = 3*10 there is only 1 minus sign, and it stays you get 6x -9y = 30

OpenStudy (anonymous):

Oh, okay. I get it. Sorry, I got kicked off for a few minutes.

OpenStudy (anonymous):

Then, you add them and the 6's cancel out right? Then you would have 0x+17y=34 right or would it be a -?

OpenStudy (anonymous):

Okay, thank you so much for your help! I finally figured out the answer.

OpenStudy (phi):

now add the 2 equations -6x+ -8y=4 6x+ -9y=30